Don Lavigne, an office equipment sales representative, earns a weekly salary plus a commission on his sales. One week his total compensation on sales of $4000 was $660. The next week his total compensation on sales of $6000 was $740. Find Don's weekly salary and commission rate.
Answer by Theo(13342)

his total compensation is equal to his salary plus his commission.

his commission is equal to his total sales times his commission rate.

in one week he earned 660.
in the next week he earned 740.

let w = his total weekly compensation.
let p = his weekly salary.
let s = his weekly sales.
let r = his commission rate.

the required formula is w = p + s * r

when w = 660 and s = 4000, this formula becomes 660 = p + 4000 * r

when w = 740 and s = 6000, this formula becomes 740 = p + 6000 * r

these are 2 equations that need to be solved simultaneously.

subtract the first equation from the second to get 80 = 2000 * r

solve for 4 to get r = 80 / 2000 = .04

the formula of 660 = p + 4000 * r now becomes 660 = p + 4000 * .04.

the formula of 740 = p + 6000 * r now becomes 740 = p + 6000 * .04.

simplify these formulas go get:

660 = p + 160

740 = p + 240

solve for p in each of these formulas to get p = 500 in both formulas.

the first week, his total compensation is 500 + .04 * 4000 = 660.

in the second week, his total compensation is 500 + .04 * 6000 = 740.

your solution is that his weekly salary is equal to 500 and his commission rate is equal to 4%.